### v1

# More Tokens, Lower Precision: Towards the Optimal Token-Precision Trade-off in KV Cache Compression

Abstract
As large language models (LLMs) process increasing context windows, the memory usage of KV cache has become a critical bottleneck during inference. The mainstream KV compression methods, including KV pruning and KV quantization, primarily focus on either token or precision dimension and seldom explore the efficiency of their combination. In this paper, we comprehensively investigate the token-precision trade-off in KV cache compression. Experiments demonstrate that storing more tokens in the KV cache with lower precision, i.e., quantized pruning, can significantly enhance the long-context performance of LLMs. Furthermore, in-depth analysis regarding token-precision trade-off from a series of key aspects exhibit that, quantized pruning achieves substantial improvements in retrieval-related tasks and consistently performs well across varying input lengths. Moreover, quantized pruning demonstrates notable stability across different KV pruning methods, quantization strategies, and model scales. These findings provide valuable insights into the token-precision trade-off in KV cache compression. We plan to release our code in the near future.
More Tokens, Lower Precision: Towards the Optimal Token-Precision
Trade-off in KV Cache Compression
Jiebin Zhang   Dawei Zhu   Yifan Song   Wenhao Wu  Chuqiao Kuang   Xiaoguang Li   Lifeng Shang   Qun Liu   Sujian Li   Peking University  Huawei Technologies zhangjiebin@pku.edu.cn
1 Introduction
As large language models (LLMs) have been widely used in various scenarios, such as document summarization (Zou et al., 2024), code completion (Roziere et al., 2023) and agent framework (Shridhar et al., 2020), there is a growing demand for models with larger context windows to handle more extensive inputs. As a result, GPT-4 (Achiam et al., 2023) have been extended to support 200k input tokens, and Gemini 1.5 (Reid et al., 2024) to 10M tokens. However, these powerful long-context capabilities come at the expense of significantly increased memory storage for the cached key and value (KV) states  (Waddington et al., 2013). Take Llama3-8B (Dubey et al., 2024) as an example, storing the KV cache of a single message with 100k tokens requires a high memory overhead of 20GB. Furthermore, as the decoding process heavily relies on the GPU memory bandwidth, the extensive KV cache also leads to dramatically increased decoding time (Fu, 2024).
To efficiently serve LLMs, various approaches have been proposed to compress KV cache (Pope et al., 2023). The predominant methods involve compressing the KV cache along two primary dimensions: token or precision. For the token dimension, KV pruning (or KV eviction) methods eliminate unimportant tokens to maintain a fixed KV cache size (Xiao et al., 2023; Zhang et al., 2024b; Ren and Zhu, 2024; Li et al., 2024). For the precision dimension, KV quantization technique reduces memory usage by approximating KV cache with lower precisions, like 8-bit or even lower (Sheng et al., 2023; Liu et al., 2024c; Hooper et al., 2024; Yang et al., 2024b). However, these existing works focuses only one dimension, either token or precision, leaving the trade-off between these two orthogonal factors largely under-explored.
Figure 1: More tokens, lower precision in KV cache leads to better performance.
In this paper, we comprehensively investigate the token-precision trade-off in KV cache compression. First, we examine the feasibility of combining KV pruning and quantization under fixed budget. We demonstrate that storing more tokens in the KV cache with lower precision, which we call quantized pruning, can significantly enhance the long-context performance of LLMs. For example, with the same KV cache budget, storing 4×4$\times$4 × tokens in 4-bit precision outperforms storing 1×1$\times$1 × tokens in 16-bit precision across various downstream long-context tasks with various input length, as shown in Figure 1. Moreover, in extremely low-resource scenarios, quantized pruning effectively preserves performance, whereas relying solely on KV pruning or quantization often leads to a significant performance collapse.
Furthermore, we conduct in-depth analysis regarding token-precision trade-off from series of key aspects, including the impact on various downstream task types and input lengths, model scaling effect, ablation on quantization strategies and fine-grained exploration of layer-wise quantized pruning. From the extensive experiments, we observe that quantized pruning achieves substantial improvements in retrieval-related tasks and consistently performs well across varying input lengths. Moreover, quantized pruning demonstrates strong feasibility across different KV pruning methods, quantization strategies, and model scales, showcasing notable stability. The analysis on token-precision trade-off presents a more substantial compression potential compared to compressing along a single dimension. We believe that our findings could offer valuable insights for developing more effective KV compression strategies in future research.
2 Related Work
KV Pruning
KV pruning compresses the KV cache along the token dimension by selectively removing unimportant tokens to reduce memory usage. Mainstream methods typically identify important tokens based on attention scores, as seen in (Liu et al., 2024b; Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024). Other methods use alternative factors such as initial tokens (Xiao et al., 2023), variance (Ren and Zhu, 2024), special tokens (Ge et al., 2024) or the L2 norm (Devoto et al., 2024) to determine token importance. Recent studies delve deeper into optimizing the allocation of KV cache memory budgets. Some explore KV cache budget allocation strategies across layers (Cai et al., 2024; Yang et al., 2024a), while other studies explore head-level KV cache budget allocation (Feng et al., 2024; Tang et al., 2024; Fu et al., 2024; Xiao et al., 2024).
KV Quantization
KV quantization compress KV cache from the precision dimension by storing KV cache using a reduced number of bits. FlexGen (Sheng et al., 2023) utilizes group-wise 4-bit quantization for both key and value cache. KIVI (Liu et al., 2024c) applies per-channel quantization on key cache and per-token quantization on value cache. KVQuant (Hooper et al., 2024) and CQ (Zhang et al., 2024a) use RoPE-related quantization, while KVQuant also preverses outliers without quantization. Atom (Zhao et al., 2024) reorders the outlier channels for fine-grained group quantization with mixed-precision. GEAR (Kang et al., 2024) uses low-rank approximation for quantization. QAQ (Dong et al., 2024) and MiKV (Yang et al., 2024b), inspired by the KV pruning methods, store discarded tokens using lower bit precision while retaining important tokens in full precision.
Other KV Compression Methods
Compressing KV cache from other dimensions typically requires modifying the model architecture, which usually necessitates additional training for adaptation. For the layer dimension, LCKV (Wu and Tu, 2024), CLA (Brandon et al., 2024) and MLKV (Zuhri et al., 2024) reduce memory usage by sharing the KV cache across adjacent layers. ShortGPT (Men et al., 2024) and DynamicSlicing (Dumitru et al., 2024) achieve compression by eliminating redundant layers. YOCO (Sun et al., 2024) changes the model structure and shares a single global KV cache across layers. For the head dimension, MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023) share the KV cache within each head groups. DeepSeek-v2 (Liu et al., 2024a) employs dimension-reduction techniques to compress all heads into a single low-rank vector. These lines of work is orthogonal to ours, as they can also be combined together.
3 Preliminaries
The decoder-only transformer model consists of a stack of transformer decoder blocks, each comprising two main components: self-attention module and the feed-forward network (FFN) module. During inference, KV cache is implemented within the self-attention module and operates in two distinct phases: i) the prefill phase, where the input prompt is used to generate KV cache for each transformer layer of LLMs; and ii) the decoding phase, where the model uses KV cache to generate the next token, and updates the KV cache with the new token.
Prefill Phase.
Let $\bm{X}\in\mathbb{R}^{b\times l_{\text{prompt}}\times d}$ be the input tensor, where b is the batch size, $l_{prompt}$ is the length of the input prompt, and d is the model hidden size. For clarity, we omit the layer index here. The key and value tensors can be computed by:
$\bm{X_{K}}=\bm{XW_{K}},\bm{X_{V}}=\bm{XW_{V}}$ (1)
where $\bm{W_{K}},\bm{W_{V}}\in\mathbb{R}^{d\times d}$ are the key and value layer weight. $\bm{X_{K}},\bm{X_{V}}$ are cached in the memory for utilization in the subsequent decoding phase.
Decoding Phase.
Let $\bm{h}\in\mathbb{R}^{b\times 1\times d}$ be hidden state of current input token. $\bm{h_{K}}=\bm{hW_{K}}$ and $\bm{h_{V}}=\bm{hW_{V}}$ are the current key and value states. $\bm{h_{K}}$ and $\bm{h_{V}}$ are first employed to update the KV cache:
$\bm{X_{K}}$ $\xleftarrow{}\text{Concat}(\bm{X_{K},h_{K}}),$
$\bm{X_{V}}$ $\xleftarrow{}\text{Concat}(\bm{X_{V},h_{V}})$ (2)
then attention output $\bm{h_{O}}$ is calculated by:
$\bm{h_{O}}=\text{Softmax}(\bm{h_{Q}X_{k}^{T}})\bm{X_{V}}$ (3)
where $\bm{h_{Q}}=\bm{hW_{Q}}$ is the output of the query layer. For ease of illustration, we ignore the FFN module and other parts of the inference workflow that are not addressed in our approach.
KV Quantization
The B-bit KV quantization process during the prefill phase can be expressed as follows: First, determine the minimum number $z_{i}$ and the maximum number $m_{i}$ in $\bm{G}_{i}$, where $\bm{G}_{i}$ is a group of number in $\bm{X}_{K}$ or $\bm{X}_{V}$. Using these numbers, compute the quantized result Q($\bm{G}_{i}$) for each group according to the formula:
$Q(\bm{G}_{i})$ $=\left\lfloor\frac{\bm{G}_{i}-z_{i}}{s_{i}}\right\rceil,\quad$ $s_{i}=\frac{m_{i}-z_{i}}{2^{B}-1}$ (4)
[…]
$\bm{G_{i}}^{\prime}=Q(\bm{G_{i}})\cdot s_{x}+z_{x}$ (5)
Pruning Method LongBench
Token=128 Token=512 Token=2048
16-bit 8-bit 4-bit 2-bit 16-bit 8-bit 4-bit 2-bit 16-bit 8-bit 4-bit 2-bit
StreamingLLM 32.1 32.2 31.7 19.1 34.6 34.5 33.9 20.7 38.1 38.2 37.8 23.8
H2O 35.6 35.6 34.7 15.8 37.5 37.4 36.7 17.7 39.8 39.7 39.0 21.1
SnapKV 35.7 35.7 35.1 16.6 40.3 40.4 39.7 20.2 41.7 41.7 41.0 22.9
PyramidKV 37.4 37.3 36.4 17.5 40.3 40.3 39.6 20.9 41.8 41.8 41.3 23.6
Pruning Method Needle-in-a-Haystack
Token=128 Token=512 Token=2048
16-bit 8-bit 4-bit 2-bit 16-bit 8-bit 4-bit 2-bit 16-bit 8-bit 4-bit 2-bit
StreamingLLM 27.7 27.7 27.5 30.9 35.3 35.3 35.5 37.3 66.4 66.5 66.4 61.8
H2O 46.9 46.6 46.8 36.4 91.2 91.1 91.0 54.8 100 100 100 74.1
SnapKV 83.7 83.7 82.5 55.9 97.4 97.4 97.2 66.3 100 100 100 78.1
PyramidKV 98.9 98.9 98.8 67.5 100 100 100 78.6 100 100 100 79.6
Table 1: Feasibility of quantized pruned tokens on LongBench and Needle-in-a-Haystack with Llama-3-8B-Instruct as backbone model. We use four KV pruning methods to retain 128, 512 and 2048 tokens, and report the results of further quantization.
Figure 2: The token-precision trade-off under varying memory budgets on LongBench and NIAH. We report the results of SnapKV-based and PyramidKV-based quantized pruning on Llama-3 and Mistral-v0.2.
KV Pruning
The goal of KV pruning is to find two submatrices $\bm{X}_{K}^{e},\bm{X}_{V}^{e}\in\mathbb{R}^{b\times s\times d}$ from the full matrices $\bm{X}_{K}$ and $\bm{X}_{V}$ during the prefill phase, given a cache budget $s<n$, while maximizing performance preservation. During the decoding phase, LLMs with KV pruning only use $\bm{X}_{K}^{e}$ and $\bm{X}_{V}^{e}$ to update KV cache and generate new tokens.
$\bm{X}_{K}^{e}$ $\xleftarrow{}\text{Concat}(\bm{X_{K}^{e},h_{K}}),$
$\bm{X}_{V}^{e}$ $\xleftarrow{}\text{Concat}(\bm{X_{V}^{e},h_{V}})$ (6)
Quantized Pruning
Quantized pruning uses KV pruning methods to obtain $\bm{X}_{K}^{e}$ and $\bm{X}_{V}^{e}$ first, and then quantizes the preserved KV states $\bm{X}_{K}^{e}$ and $\bm{X}_{V}^{e}$ to Q($\bm{X}_{K}^{e}$) and Q($\bm{X}_{V}^{e}$) using various KV quantization methods in the prefill phase. In the decoding phase, the dequantized results from Q($\bm{X}_{K}^{e}$) and Q($\bm{X}_{V}^{e}$) are used to generate new tokens.
4 Experimental Setup
Benchmarks
We assess the performances of quantized pruning using LongBench (Bai et al., 2024) dataset and Needle-in-a-Haystack (Kamradt, 2023) test. We also employ RULER (Hsieh et al., 2024), a dataset with different input length and diverse types of needles across 4 task categories, to better access the impact of input length in Section 6.1. More details can be seen in Appendix A.
LLMs
We use state-of-the-art open-weight LLMs, including Llama-3-8B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023). For scaling experiments in Section 6.2, we also test the performance of Llama-3.2-1B and Llama-3.2-3B (Dubey et al., 2024).
Setup
We aim to comprehensively investigate the token-precision trade-off in KV cache compression. We report the ratio of compressed KV cache and the full KV cache for memory budge. For KV pruning, we employ PyramidKV (Cai et al., 2024) and SnapKV (Li et al., 2024), recognized as leading methods in diverse scenarios. We also include H2O (Zhang et al., 2024b) and StreamingLLM (Xiao et al., 2023) to evaluate the feasibility of quantized pruning in Section 5. For KV quantization, we adopt KIVI (Liu et al., 2024c) as the default method due to its stability and broad compatibility. Moreover, in Section 6.3, we examine the effects of quantization strategies from FlexGen (Sheng et al., 2023) and KVQuant (Hooper et al., 2024) for a comprehensive comparison. We use HQQ (Badri and Shaji, 2023) framework to perform quantization on KV cache. More details can be seen in Appendix B.
5 Optimal Token-Precision Trade-Off
In this section, we aim to find the optimal token-precision trade-off in KV cache compression. We frist examine the feasibility of combining KV pruning and quantization(Q1). Subsequently, we explore the best optimal allocation strategy between precision and token under varying memory budgets(Q2).
{bclogo}
[…]
We first evaluate the feasibility of quantized pruned KV cache as a prerequisite for exploring the token-precision trade-off. We use Llama-3-8B-Instruct and evaluate various KV pruning methods on the LongBench and NIAH. We report the results of quantizing the remaining tokens to different precision levels after applying KV pruning.
From Table 1, we observe that it is feasible to quantize pruned KV cache for a lowe compression rate.
For most KV pruning methods we evaluate, further quantizing the preserved tokens to as low as 4-bit precision results in minimal performance degradation, but quantizing to 8-bit precision shows negligible impact. However, reducing precision to 2-bit leads to a drastic performance decline across most KV pruning methods.
This observation holds consistently across different KV pruning methods and varying numbers of preserved tokens.
Compared with precision, reducing the number of preserved tokens leads to more significant performance degradation. Specifically, when the number of preserved tokens is reduced to 1/4 (from 2048 to 512), all KV pruning methods experience a noticeable performance drop. In contrast, when the precision is reduced to 1/4 (from 16-bit to 4-bit), which has the same memory budget as token dimension, the performance degradation is relatively mild. This suggests that, under the same memory budget, tokens might have a more significant impact on the results compared to precision.
{bclogo}
[…]
Observing that KV pruning and KV quantization can be effectively combined, we further investigate that, given a fixed memory budget, how to balance the trade-off between number of preserved tokens and precision to achieve optimal performance. To this end, we evaluate the performance of quantized pruning using two leading KV pruning methods, SnapKV and PyramidKV, across different memory budgets on LongBench and NIAH.
Specifically, we compare three configurations with approximately equivalent memory usage: 1) Using standalone KV pruning to retain 1×1$\times$1 × tokens in 16-bit precision. 2) Quantized pruning by retaining 2×2$\times$2 × tokens in 8-bit precision. 3) Quantized pruning by retaining 4×4$\times$4 × tokens in 4-bit precision.
As shown in Figure 2, we observe that quantized pruning, which preserves more tokens at a lower precision, consistently outperforms standalone KV pruning methods across various budgets. For the NIAH task, the improvements from quantized pruning are particularly pronounced. This may be attributed to that quantized pruning can cover more tokens for retrieval under the same memory budget compared to standalone KV pruning.
In high-budget scenarios, the 8-bit strategy tends to deliver slightly better performance, which may due to the number of tokens at this budget is already quite large. In low-budget scenarios, such as 1/128 KV cache budget, storing more tokens at 4-bit precision yields superior results, highlighting the importance of token coverage when resources are constrained. Overall, using lower precision to preserve more tokens under a limited budget results in notable performance gains, compared to standalone KV pruning methods that use full precision to store fewer tokens.
Summary
We demonstrate that storing more tokens in the KV cache with lower precision can significantly enhance the long-context performance of LLMs under fixed KV cache memory budget.
6 Further Analysis
In this section, we further investigate series of key aspects regarding token-precision trade-off, including the impact of quantized pruning on various downstream task types and input lengths, model scaling effect, ablation on quantization strategies and fine-grained exploration of layer-wise quantized pruning.
Models Token Bit Task Types
SQA MQA SUMM Fewshot Syn. Code RULER-8k
Llama-3-8B-Instruct 512 16 28.2 31.9 23.5 67.6 37.7 57.6 67.5
1024 8 29.6 33.1 24.3 67.9 37.4 58 74.9
2048 4 30.7 32.5 25.3 68.8 37.2 57.6 82.2
Mistral-7B-Instruct-v0.2 512 16 33.7 27.3 24.3 65.6 41.75 54 53.1
1024 8 34.2 29 25.6 66.4 43.73 54.8 62.1
2048 4 35.2 28.14 26.6 66.9 43.08 55.4 73.6
Table 2: The token-precision trade-off in different task types. We report the results of 6 task types in LongBench and 8k subset of RULER. We use PyramidKV-based quantized pruning.
6.1 Impact on Downstream Tasks
Task Types
To further investigate the token-precision trade-off in different task types, we evaluate PyramidKV-based quantized pruning on six task types from LongBench and the 8K subset of the RULER dataset. We use PyramidKV with 512 retained tokens as the baseline, and explore the token-precision trade-off under this fixed memory budget, as this setting exhibits minimal performance differences across three precision levels, making it easier to assess the impact of task types.
As illustrated in Table 2, we observe that the performance of quantized pruning is remarkably consistent across different task types. Specifically, lower precision, which retains more tokens in KV cache, leads to substantial performance improvements in the RULER task, which heavily relies on retrieving content from the input. Tasks with high retrieval demands, such as Summarization and Single-Doc QA, also show noticeable gains with quantized pruning, particularly when 4×4$\times$4 × tokens are preserved at 4-bit precision.
For tasks requiring more reasoning rather than intensive retrieval, such as Code Completion, Synthetic and Multi-Doc QA, the benefits of trading precision for more tokens are less pronounced. In these cases, storing fewer tokens with higher precision generally performs better. For example, using 1024 tokens in 8-bit precision achieves the hightest score of 58 in Code task with Llama-3.
Input Lengths
Figure 3: The token-precision trade-off in different input lengths. We report the results of LongBench and three subsets of RULER. We use PyramidKV-based quantized pruning.
To evaluate the token-precision trade-off across various input lengths, we conduct experiments on subsets with different input length of the RULER dataset. Additionally, we analyze LongBench by grouping its data based on input length. The results is shown in Figure 3 and more detailed information can be found in Appendix A.
Our observations are as follows: quantized pruning consistently outperforms standalone KV eviction methods across various input length, regardless of the models and task types. Within the same dataset, scores decrease as input length increases; however, the relative differences among different compression methods remain similar across varying input lengths. Moreover, quantized pruning achieves significant performance improvements across all input lengths for retrieval demanded tasks like RULER.
6.2 Scaling Effect on Quantized Pruning
Figure 4: Scaling effect on Llama family models, with PyramidKV-based quantized pruning. All models are under fixed ratio of KV cache budget.
To investigate the impact of model scaling on quantized pruning, we conducted experiments on three models from the Llama series: Llama3-8B, Llama3.2-3B, and Llama3.2-1B. For both the Base models and Instruct models, we evaluated their performance on LongBench under two fixed KV cache budgets: 1/16 and 1/64.
As shown in Figure 4, we observe that quantized pruning consistently achieves better performance across all scaling levels. The performance gap between quantized pruning and standalone KV pruning methods remains relatively stable across different model scales. Notably, when the KV cache budget is relative small to 1/64 , the performance improvement brought by quantized pruning is higher compared to 1/16 KV cache budget, which aligns with the conclusions we observed earlier in Q2. For Base models, although the performance improvement from scaling is smaller compared to Instruct models, quantized pruning still provides a noticeable performance boost.
These findings highlighting the robustness and effectiveness of quantized pruning across model scaling.
6.3 Ablation on Quantization Strategies
Figure 5: Ablation of quantization strategies on quantized pruning, remaining 512 KV tokens in 4-bit.
While there has been extensive research on strategies for KV cache quantization, it remains unclear whether existing quantization strategies remain effective when combined with KV pruning methods. In this section, we aim to investigate the impact of KV quantization strategies and group size on quantized pruning, and present our results in Figure 5 and Table 3.
Quantization methods
We explore the methods in FlexGen (Sheng et al., 2023), KIVI (Liu et al., 2024c), and KVQuant (Hooper et al., 2024). To elaborate, for the FlexGen methods, KV quantization is applied to both the key and value caches along the token dimension, grouping every 64 elements without filtering outlier numbers. We modify the FlexGen by (1) filtering 1% of outlier numbers in both the key and value caches, as mentioned in KVQuant (2) quantizing the key along the channel dimension, as in KIVI and (3) combining (1) and (2). These correspond to the results labeled as FlexGen+Outlier 1%, KIVI, and KIVI+Outlier 1% in the Figure 5.
We can observe that none of the quantization strategies show significant performance degradation when combined with KV pruning methods, demonstrating the relative stability of quantized pruning. The KIVI method consistently outperforms FlexGen across various models and KV pruning methods. The improvement is particularly pronounced for PyramidKV on the Mistral model, underscoring the significance of quantizing key states along the channel dimension. Filtering 1% of outlier numbers proves effective for the FlexGen strategy but yields limited improvements for KIVI. It shows some benefit on Llama models but can result in negative gains on the Mistral model.
Overall, KIVI demonstrates strong performance when combined with KV pruning methods, while other KV quantization strategies also maintain good results, highlighting the stability of quantized pruning.
Group Size
We then analyzed the impact of group size during KV quantization. Using the SnapKV and PyramidKV methods to retain 512 tokens, we experimented with 4-bit quantization and observed the performance variations when the group sizes were set to 32, 64 and 128.
Model Method GroupSize
32 64 128
Llama-3 SnapKV 40.42 39.55 38.87
PyramidKV 40.33 39.65 38.91
Mistral-v0.2 SnapKV 40.45 40.30 40.09
PyramidKV 40.31 40.49 40.03
Table 3: The impact of group size for quantized pruning on LongBench, remaining 512 KV tokens in 4-bit.
As shown in Table 3, smaller group sizes lead to performance improvements at the cost of higher memory usage. Reducing the group size from 128 to 64 resulted in a notable improvement, but further decreasing it from 64 to 32 yielded minimal gains for the Mistral model. Therefore, we set the default quantization group size to 64 to balance performance and memory usage in our experiments.
6.4 Exploration on Layer-Wise Quantized Pruning
Inspired by the observation that different layers may have varying requirements for the number of tokens in PyramidKV (Cai et al., 2024) and PyramidInfer (Yang et al., 2024a), we further investigate whether the demands for precision and preserved tokens are consistent across layers. To explore this, we use the best-performing configuration from previous experiments, 4-bit precision with 4×4$\times$4 × tokens, as the baseline and compare it against layer-wise configurations adopting 8-bit precision with 2×2$\times$2 × tokens and 16-bit precision with 1×1$\times$1 × tokens. Using SnapKV as the KV pruning method, we present the results for Llama-3 and Mistral-v0.2 under two budget constraints in the Figure 6. Configurations are modified every 4 layers for the initial and final layers, while intermediate layers are reconfigured every 8 layers. The x-axis indicates the layers where the modified configurations are applied, while the y-axis shows the relative change to the baseline (4-bit precision with 4×4$\times$4 × tokens) on LongBench and RULER-4k.
Figure 6: The results of layer-wise quantized pruning on Llama-3-8B-Instruct, with SnapKV as pruning method. We use 4×4$\times$4 × KV token 4-bit as baseline and report the relative change.
Initially, it is evident that for most layers, transitioning from 4×4$\times$4 × tokens with 4-bit precision to higher precision and fewer tokens results in a performance decline under constrained KV cache budgets. Specifically, the shift to 8-bit shows a relatively minor performance drop, whereas moving to 16-bit with fewer preserved tokens leads to a more significant decrease. These layers-wise trade-off conclusions are consistent with our experiments before.
Notably, modifying intermediate layers causes a drastic performance decline, while adjustments made at the initial and final layers result in comparatively smaller performance reductions. This effect is especially pronounced in retrieval-related tasks such as RULER-4k, where significant performance differences are observed. On LongBench, changes are less significant, with a notable performance drop only observed at 16-bit precision. These findings highlight that, under the same memory budget, preserving more tokens in intermediate layers is crucial for the performance, while the token-precision trade-off in the initial and final layers exerts a more balanced influence on the results.
7 Conclusion
We investigate a series of critical yet unexplored questions regarding the effectiveness and feasibility of token-precision trade-off in KV cache compression. Through comprehensive experiments, we demonstrate that storing more tokens in the KV cache with lower precisioncan significantly enhance the long-context performance of LLMs, and demonstrating robust performance across diverse input lengths, downstream tasks, with particularly significant gains in retrieval tasks. Moreover, we find quantized pruning demonstrates strong feasibility across different KV pruning methods, quantization strategies, and model scales. Our analysis sheds light on the token-precision trade-off of KV cache memory optimization, offering valuable insights into designing more efficient compression strategies. We hope this work deepens our understanding of KV cache compression and inspires future research.
Limitations
While our work demonstrates the effectiveness of KV compression through trade-offs between token and precision dimensions, other potential dimensions, such as head and layer, remain unexplored. Investigating the feasibility of combining these dimensions with token and precision for a more substantial compression potential represents an avenue for future research. Additionally, the current implementation of quantized pruning suffers from inefficiencies in dequantizing the KV cache, hindering the full realization of speedup benefits from the memory savings. In future work, we aim to address this issue by optimizing the implementation, such as integrating fusion operators to combine dequantization with matrix multiplication.
Appendix A Datasets
LongBench
LongBench (Bai et al., 2024) includes 17 datasets covering 6 categories of tasks, which can be divided into single-document QA (Dasigi et al., 2021; Kočiskỳ et al., 2018), multi-document QA (Yang et al., 2018; Ho et al., 2020), summarization (Huang et al., 2021; Fabbri et al., 2019; Zhong et al., 2021), few-shot learning (Gliwa et al., 2019; Joshi et al., 2017; Li and Roth, 2002), synthetic, and code generation (Guo et al., 2023; Liu et al., 2023). LongBench features an average input length ranging from 1,235 to 18,409 tokens. For inputs exceeding the model’s context window length, we split the data and only take the beginning and end segments of the input to fill the context window length. For Input Lengths Experiment, we select datasets with sufficient data to cover three input length ranges: (<4k, 4k 8k, and >8k). These datasets include MultiFieldQA-en, 2WikiMultihopQA, GovReport, TREC, TriviaQA, SAMSum, and RepoBench-P, representing a variety of task types. We refer to the three subsets as LB-4k, LB-8k, and LB-16k, respectively.
NIAH
Needle-in-a-Haystack(NIAH) (Kamradt, 2023) is a challenging pressure test designed to assess the ability of models to accurate identify and retrieve relevant information from lengthy context. NIAH randomly inserts key information into an arbitrary position within a long essay. In our setup, we use PaulGrahamEssays as the haystack and the sentence "The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day." as the needle, which is the default setting of NIAH. We vary the essay length from 1,000 tokens up to the models’ context window limits, increasing by 100 tokens per step for Llama-series models and 400 tokens per step for Mistral.
RULER
RULER (Hsieh et al., 2024) generates synthetic examples to evaluate long-context language models with configurable sequence lengths and varying task complexities. It includes four task categories: Retrieval, Multi-hop Tracing, Aggregation, and Question Answering. The dataset comprises six subsets with input lengths of 4K, 8K, 16K, 32K, 64K and 128K tokens. In our experiments, we use the 4K, 8K and 16K subsets to test the models within their context window limits.
Figure 7: The results of Layer-Wise Quantized Pruning on Mistral-7B-v0.2-Instruct.
Appendix B Experiment Setup
Memory Budgets
We report the ratio of compressed KV cache and the full KV cache for memory budge. The full KV cache for Llama-3 is 8k KV tokens in 16-bit on both LongBench and NIAH, while for Mistral-v0.2 is 16k KV tokens on LongBench and 32k KV tokens on NIAH in 16-bit.
KV eviction methods
We retain the last 32 tokens for StreamingLLM (Xiao et al., 2023), H2O (Zhang et al., 2024b), and SnapKV (Li et al., 2024), while keeping 8 tokens for PyramidKV (Cai et al., 2024), as recommended in the corresponding paper (Cai et al., 2024). For other settings, we adopt the default configurations from the PyramidKV codebase.
KV quantization
We utilize HQQQuantizedCache from Huggingface and adjust the group dimensions of keys and values to implement grouped quantization strategies from FlexGen (Sheng et al., 2023) and KIVI (Liu et al., 2024c). We use 64 as the default group size which is suggested in FlexGen (Sheng et al., 2023).In the experiments involving outlier filtering, we exclude numbers in the KV cache with a absolute value exceeding 6 from quantization, which roughly corresponds to the top 1% of outliers based on our validation set analysis.
Appendix C More results in Experiments
Layer-Wise Quantized Pruning
We alse report the results for Mistral-v0.2 in Figure 7, the results are similiar to Llama-3.